OpenStudy (anonymous):

dy/dx=6sin^2x cosx, y(0)=6. Solve the initial value problem.

OpenStudy (cruffo):

Looks like you could separate variables, and some u-substitution. Familiar with those ideas?

OpenStudy (anonymous):

Yes I am familiar with substituting u to find the answer.

OpenStudy (cruffo):

Great! So lets start by seperating variables: \[\frac{dy}{dx} = \sin^2(x)\cos(x)\] \[1dy = \sin^2(x)\cos(x)dx\] \[\int 1dy = \int \sin^2(x)\cos(x)dx\]

OpenStudy (cruffo):

let \(u = \sin(x)\) the \(du = \cos(x) dx\)

OpenStudy (anonymous):

Alright, so I see that you started out with the chain rule and then you substituted u

OpenStudy (cruffo):

Yah.... chain rule :)

OpenStudy (anonymous):

Alright, I'm following you. You got 3sin^2x from dividing 6/2. So does sin^2x turn into (1-cos^2x)?

OpenStudy (cruffo):

I wasn't leaning toward using any identity, but you could.... Note: \[\int 1\; dy = \int \sin^2(x) \cos(x) dx = \int u^2du\] by substitution. So just power rule.

OpenStudy (anonymous):

Okay

OpenStudy (cruffo):

The left hand side \(\int 1 dy = y\) The right hand side \(\int u^2 du = \dfrac{1}{3} u^3 + c = \dfrac{1}{3}\sin^3(x) + c\)

OpenStudy (cruffo):

\[y = \frac{1}{3}\sin^3(x) + c\] The rest of the problem is figuring out what c is.